# HumanGif: Single-View Human Diffusion with Generative Prior

### v2

@@ 4.1 Experimental Setup
Datasets. We evaluate the performance of our HumanGif on two multi-view human modelling datasets, i.e., RenderPeople [1] (Synthetic Humans) and DNA-Rendering [22] (Real-World Humans). For RenderPeople, we randomly sample 450 subjects as the training set and 30 subjects for testing. For each subject, we use all frames from the training data for training, 6 frames with 4 camera viewpoints for novel view synthesis, and all frames with a front camera view for novel pose synthesis. For the DNA-Rendering dataset, we use 416 sequences from Part 1 and 2 for training and 10 sequences for evaluation. For each subject, we use all frames from the training data for training, 48 frames with 4 camera viewpoints for novel view synthesis, and all frames with a front camera view for novel pose synthesis. The foreground masks, camera, and SMPL parameters from these two datasets are used for evaluation purposes.
Comparison Methods. We compare our HumanGif with two categories of state-of-the-art single-view-based animatable human modelling methods, i.e., a generalizable Human NeRF method, SHERF [49], and diffusion-based 2D character animation methods, AnimateAnyone [47] and Champ [143]. For fair comparisons, we retrain SHERF and Champ by using their official codebase and AnimateAnyone with the open-source implementation from MooreThreads [^1] https://github.com/MooreThreads/Moore-AnimateAnyone.
Implementation Details. The training phase is divided into three stages. In the first stage, we process all frames by resizing and cropping the images to be 512x512 resolution for the RenderPeople dataset and 768x768 resolution for the DNA-Rendering dataset. Then we train a model to learn the mapping from a reference image and a target pose image to a target human image. This stage is trained for 150,000 iterations with a batch size of 4. In the second stage, we incorporate the multi-view attention layer after the spatial-attention and cross-attention in the denoising UNet to learn multi-view consistency. This stage is trained for 10,000 iterations with a batch size of 24 frames (a video length of 6 and 4 views) on each GPU. In the last training stage, we train the temporal layer with the same batch size used in the second stage for 10,000 iterations.
Figure 3: Qualitative results of novel view synthesis (1st and 3rd row) and novel pose synthesis (2nd and 4th row) produced by SHERF, Animate Anyone, Champ and our HumanGif on RenderPeople and DNA-Rendering datasets.
@@ 4.2 Quantitative Results
As shown in Tab. 1, HumanGif outperforms baseline methods in perceptual quality metrics (LPIPS) across both datasets. While SHERF, the state-of-the-art generalizable animatable Human NeRF method for single-view image input, achieves the highest PSNR scores in both novel view and pose synthesis tasks, it often produces blurred images, particularly for unseen views and poses. This is attributed to its limited capability to infer missing information from a single input image, which explains its subpar performance in perceptual quality metrics (LPIPS). Animate Anyone and Champ, 2D SOTA character animation methods, demonstrate reasonably good perceptual quality in unseen views and poses by leveraging the generative prior of Stable Diffusion. However, they fail to capture detailed information from the input image. Meanwhile, these methods fail to produce consistent novel view and pose results. In contrast, our HumanGif leverages generative prior, a Human NeRF module, and a multi-view and temporal attention mechanism, enabling it to effectively learn fine-grained details from the input image and produce consistent novel view and pose results, even in the single-image setting. Beyond image quality metrics, we also evaluate video fidelity for animatable human videos generated by these methods. Our HumanGif achieves the best FVD metric with a temporal-attention module, demonstrating superior temporal consistency. At the same time, SHERF performs reasonably well in video fidelity due to its effective LBS modeling. We incorporate a user study in Fig. 4 to show human perceptual results.
Table 2: Ablation study on DNA-Rendering. The left side shows different design components that are ablated on.
Generative Prior Human NeRF Module Image-Level Loss Multi-View Attention Motion Attention Novel View Novel Pose
@@
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 23.647 0.935 0.049 23.504 0.931 0.048 11.24
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 23.686 0.935 0.047 24.275 0.935 0.045 9.88
Figure 4: User preference scores on synthesis results.
@@ 4.3 Qualitative Results
We show qualitative results of novel view synthesis (1st and 3rd row) and novel pose synthesis (2nd and 4th row) of our HumanGif and baseline methods in Fig. 3. SHERF produces reasonable RGB renderings for the part visible from the input image, but it struggles to get realistic results for the part invisible from the input image. For example, it produces blurry images in the back view when given a front view image. Thanks to the generative prior involved in our HumanGif, we can produce realistic outputs (e.g., back view in the 3rd row of Fig. 1) in unseen views and poses. 2D character animation methods, Animate Anyone and Champ, produce realistic results, but they fail to recover fine-grained details (e.g., patterns on cloths and eyeglasses) from the input image. By incorporating the Human NeRF module to enhance the information from the input 2D observation, our HumanGif successfully recovers the detailed information from the input image.
@@ 4.4 Ablation Study
To validate the effectiveness of the proposed components, we subsequently integrate them and evaluate their performance on the DNA-Rendering dataset. As shown in Tab. 2 and Fig. 5, training HumanGif without inheriting the generative prior (pre-trained weights) from Stable Diffusion results in distorted human images with incorrect textures. By incorporating the generative prior, HumanGif generates realistic human images while preserving the identity of the input. However, it still struggles to capture fine-grained details from the input image. For instance, high-heeled shoes are mistakenly added to bare feet, and the garment style in the back view differs from that in the front view (see Fig. 5). Incorporating the Human NeRF module enables HumanGif to learn detailed information from the input image, effectively addressing these issues. Additionally, introducing the image-level loss enhances the model’s ability to produce consistent results. For example, pants are corrected to a skirt in the generated output. The multi-view attention considered in this paper forces the model to learn view-consistent results (e.g., the hair in Fig. 5). Furthermore, the multi-view attention mechanism ensures the model learns view-consistent features, such as maintaining consistent hairstyles across views (as seen in Fig. 5). Lastly, integrating the temporal module refines the model’s ability to generate smooth novel pose sequences and resolves issues related to inconsistent human poses.
5 Discussion and Conclusion
@@
To sum up, we propose HumanGif, which reformulates the single-view-based 3D human novel view and pose synthesis as a single-view-conditioned human diffusion process, utilizing generative priors from foundational diffusion models to complement the missing information. To further produce perpetually realistic, view-consistent, and temporally coherent human avatars from a single image, we incorporate a Human NeRF module along with a view-attention, and temporal attention layer into our HumanGif. Furthermore, we introduce an image-level loss during optimization to bridge the gap between latent and image spaces in diffusion models. On two multi-view human datasets, our HumanGif achieves the best perceptual results.
@@
Limitation and Future Work. 1) While our proposed HumanGif improves the synthesis performance, there still exists inconsistency between the generated images and the input image. How to align target images with the input image and target poses remains a future direction to be explored. 2) There is still room to improve the visual quality of generated results. We observe that VAE produces distortions for facial areas in some cases. Introducing VAE tailored for human images is a promising direction to improve the quality. 3) Currently, our HumanGif is trained on existing multi-view human datasets. Scaling our HumanGif to larger human datasets remains a future direction to be explored.

### v3

@@ 4.1 Experimental Setup
Datasets. We evaluate the performance of our HumanGif on four human modelling datasets, i.e., RenderPeople [1] (multi-view image), DNA-Rendering [22] (multi-view video), THuman2.1 [137] (3D scan), and TikTok [60] (monocular video). For RenderPeople, we randomly sample 450 subjects as the training set and 30 subjects for testing. For each subject, we use all frames from the training data for training, 6 frames with 4 camera viewpoints for novel view synthesis, and all frames with a front camera view for novel pose synthesis. For the DNA-Rendering dataset, we use 416 sequences from Part 1 and 2 for training and 10 sequences for evaluation. For each subject, we use all frames from the training data for training, 48 frames with 4 camera viewpoints for novel view synthesis, and all frames with a front camera view for novel pose synthesis. The foreground masks, camera, and SMPL parameters from these two datasets are used for evaluation purposes. For the THuman2.1 dataset, we randomly select 2345 3D scans for training and 100 3D scans for evaluation. We render 24 multi-view images at a resolution of 512x512 for each scan, and evaluate the performance on all rendered views of the test set. For the TikTok dataset, we follow previous work [121] to split the training and test sets. We estimate SMPL and camera from a 2D image using 4D-Humans [34], and segmentation masks from SAM [73]. For THuman2.1 and TikTok datasets, we perform the training on both datasets.
Comparison Methods. We compare our HumanGif with two categories of state-of-the-art single-view-based animatable human modelling methods, i.e., a generalizable Human NeRF method, SHERF [51], and diffusion-based 2D character animation methods, AnimateAnyone [49], MagicAnimate [132] and Champ [149]. We evaluate the performance of MagicAnimate and Champ by using their released checkpoints. For fair comparisons, we also fine-tune SHERF and Champ by using their official codebase and AnimateAnyone with the open-source implementation from MooreThreads [^1] https://github.com/MooreThreads/Moore-AnimateAnyone. More implementation details are shown in Appendix.
Figure 3: Qualitative results of novel view synthesis (1st, 3rd, 5th, and 6th row) and novel pose synthesis (2nd and 4th row) produced by SHERF, Animate Anyone, Champ, and our HumanGif on RenderPeople, DNA-Rendering, and THuman2.1 datasets.
@@ 4.2 Quantitative Results
As shown in Tab. 1, HumanGif outperforms baseline methods in perceptual quality metrics (SSIM and LPIPS) across all datasets. While SHERF, the state-of-the-art generalizable animatable Human NeRF method for single-view image input, achieves the highest PSNR scores in both novel view and pose synthesis tasks, it tends to produce blurred images, particularly for unseen views and poses. This is attributed to its limited capability to infer missing information from a single input image, which explains its subpar performance in perceptual quality metrics (SSIM and LPIPS). To validate whether the original trained 2D SOTA character animation methods perform well in the novel-view and novel pose synthesis tasks, we evaluate the performance of MagicAnimate and Champ by using their released checkpoints. Although these methods achieve reasonable performance on novel pose synthesis, they fail to produce high-quality results on the novel-view synthesis task. To show fair comparisons, we fine-tune Animate Anyone and Champ on the same datasets as HumanGif. These two methods demonstrate reasonably good perceptual quality in unseen views and poses by leveraging the generative prior of Stable Diffusion. However, they fail to capture detailed information from the input image. Meanwhile, these methods fail to produce consistent novel view and pose results. In contrast, our HumanGif leverages proposed modules to effectively learn fine-grained details from the input image and produce consistent novel view and pose results, even in the single-image setting. Beyond image quality metrics, we also evaluate video fidelity for animatable human videos generated by these methods. Our HumanGif achieves the best FVD metric with a temporal-attention module, demonstrating superior temporal consistency. At the same time, SHERF performs reasonably well in video fidelity due to its effective LBS modeling. We incorporate a user study in Fig 5 to show human perceptual results. To further show generalizability on in-the-wild images, we evaluate HumanGif on in-the-wild human images by following the same data processing pipeline for TikTok.
Figure 4: Generalization performance of HumanGif on in-the-wild images.
Table 2: Ablation study on DNA-Rendering. The left side shows examined combinations of components that are ablated.
Generative Prior Human NeRF Module Image-Level Loss View/Motion Attention Novel View Novel Pose
@@
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 23.686 0.935 0.047 24.275 0.935 0.045 9.88
Figure 5: User preference scores.
@@ 4.3 Qualitative Results
We show qualitative results of novel view synthesis (1st, 3rd, 5th, and 6th row) and novel pose synthesis (2nd and 4th row) of our HumanGif and baseline methods in Fig. 3. SHERF produces reasonable RGB renderings for the part visible from the input image, but it struggles to get realistic results for the part invisible from the input image. For example, it produces blurry images in the back view when given a front-view image. Thanks to the generative prior involved in our HumanGif, we can produce realistic outputs (e.g., back view in the 3rd row of Fig. 1) in unseen views and poses. 2D character animation methods, Animate Anyone and Champ, produce realistic results, but they fail to recover fine-grained details (e.g., patterns on clothes and eyeglasses) from the input image. By incorporating the Human NeRF module to enhance the information from the input 2D observation, our HumanGif successfully recovers the detailed information from the input image.
@@ 4.4 Ablation Study
To validate the effectiveness of the proposed components, we subsequently integrate them and evaluate their performance on the DNA-Rendering dataset. As shown in Tab. 2 and Fig. 6, training HumanGif without inheriting the generative prior (pre-trained weights) from Stable Diffusion results in distorted human images with incorrect textures. By incorporating the generative prior, HumanGif generates realistic human images while preserving the identity of the input. However, it still struggles to capture fine-grained details from the input image. For instance, high-heeled shoes are mistakenly added to bare feet, and the garment style in the back view differs from that in the front view (see Fig. 6). Incorporating the Human NeRF module enables HumanGif to learn detailed information from the input image, effectively addressing these issues. Additionally, introducing the image-level loss enhances the model’s ability to produce consistent results. For example, pants are corrected to a skirt in the generated output. Furthermore, the unified view/temporal attention mechanism ensures that the model learns view-consistent features, such as maintaining consistent hairstyles across views (as seen in Fig. 6), and generates smooth novel pose sequences, and resolves issues related to inconsistent human poses.
Figure 7: Failure cases on synthesis results.
5 Conclusion and Discussion
@@
To sum up, we propose HumanGif, which reformulates the single-view-based 3D human novel view and pose synthesis as a single-view-conditioned human diffusion process, utilizing generative priors from foundational diffusion models to complement the missing information. To further produce perpetually realistic, view-consistent, and temporally coherent human avatars from a single image, we incorporate a spatially and temporally aligned conditional signal rendered from the Human NeRF module along with a Plücker ray representation, and a unified view/temporal attention layer into our HumanGif. Furthermore, we introduce an image-level loss during optimization to bridge the gap between latent and image spaces in diffusion models. Experiments demonstrate that our HumanGif achieves the best perceptual results.
@@
Limitation and Future Work. 1) While our proposed HumanGif improves the synthesis performance, there still exists an inconsistency between the generated images and the input image. For example, as shown in Fig. 7, it is still challenging to learn the correct geometry of fingers, and the generated images may contain additional accessories (e.g., watch). How to align target images with the input image and target poses remains a future direction to be explored. 2) There is still room to improve the visual quality of generated results, especially for facial areas. We observe that VAE produces distortions for facial areas in some cases. Introducing VAE tailored for human images is a promising direction to improve the quality. 3) Our HumanGif is trained on existing multi-view human datasets. Scaling our HumanGif to larger in-the-wild human datasets remains a future direction to be explored.
Appendix A Implementation Details
Training Details. All our experiments are conducted on 4 NVIDIA H100 GPUs. The training phase is divided into three stages. In the first stage, we process all frames by resizing and cropping the images to be 512x512 resolution for the RenderPeople, THuman2.1, and TikTok datasets and 768x768 resolution for the DNA-Rendering dataset. Then we train a model to learn the mapping from a reference image and a target pose image to a target human image. During optimization, the Human NeRF module is jointly optimized with Denoising UNet, ReferenceNet, Pose Encoder modules. This stage is trained for 120,000 iterations with a batch size of 8. In the second stage, we incorporate the view/temporal attention layer after the spatial-attention and cross-attention in the denoising UNet to learn multi-view consistency. We fine-tune the view/temporal attention layer while keeping the parameters of other modules fixed. This stage is trained for 10,000 iterations with a batch size of 24 frames on each GPU. We inject a class embedding [0] for novel view synthesis and a class embedding [1] for novel pose synthesis. We use the Adam [71] optimizer for all stages, set the initial learning rate as $1\times 10^{-5}$, and decay the learning rate with a linear scheduler.
3D-aware features. We include two forms of 3D-aware features in our human NeRF module, i.e., point-level features and pixel-aligned features. 1) Point-level Features. We first extract per-point features by projecting the SMPL vertices onto the 2D feature map of the input image. Next, we apply inverse LBS to transform the posed vertex features into the canonical space. These transformed features are then voxelized into sparse 3D volume tensors and further processed using sparse 3D convolutions [24]. From the encoded sparse 3D volume tensors, we extract point-level features $\bm{f}_{\text{point}}(\bm{x}^{c})$ for each point $\bm{x}^{c}$. With the awareness of 3D Human structure, point-level features capture local texture details in the seen area and infer textural information for unseen areas through sparse convolution. 2) Pixel-aligned Features. Due to limited SMPL mesh and voxel resolution, point-level features suffer from significant information loss, especially in areas visible from the reference image. To compensate for the information loss problem, we additionally extract pixel-aligned features by projecting 3D deformed points $\bm{x}^{c}$ into the input view. Each deformed point $\bm{x}^{c}$ is transformed into the observation space as $\bm{x}^{\text{ref}}=\text{LBS}(\bm{x}^{c};\bm{\theta}^{\text{ref}},\bm{\beta}^{\text{ref}})$ using LBS. It is then projected onto the input view, allowing us to query the pixel-aligned features. Leveraging the complementary strengths of point-level and pixel-aligned features, our Human NeRF module effectively captures fine-grained feature details in regions visible in the reference view while inferring features for regions occluded in the reference view.